nuggetomaster:

Rusty's hair grows at the rate of 1 over 4 inch per month. How many months will it take Rusty's hair to grow 5 over 8 inch? Explain your answer using words, and show your work using division of fractions

jhonyy9:

so do you know the rule of 3 simple ? using this you can write 1/4 inch grows in 1 month 5/8 inch grows in x month - so now cross multiplie and calcule the x value x = (5/8 *1)/(1/4 = (5/8)/(1/4) = ? month hope helped understandably easy

jackiejackiejack:

Welcome to Qc!